OpenStudy (bossimbacon):

what is the slope of the line represented by the equation 10x + 5y = 4 @alexanderhumboltd2

OpenStudy (bossimbacon):

@AlexandervonHumboldt2

OpenStudy (texaschic101):

10x + 5y = 4 5y = -10x + 4 y = -10/5x + 4/5 y = -2x + 4/5 In y = mx + b form the slope is in the m position...therefore, your slope is -2

OpenStudy (bossimbacon):

thank you!

OpenStudy (texaschic101):

no problem :)